OpenStudy (anonymous):

What is the slope-intercept form of an equation of a line parallel to y = 3x + 8 and passing through the point (-4, 4)? Question 7 options: y = 3x - 16 4y = 3x + 8 y = 3x - 8 y = 3x + 16

OpenStudy (aroub):

y=mx+b , m=3 , (-4,4) 4=3(-4)+b 4=-12+b 16=b => y=3x+16

OpenStudy (anonymous):

you are the smartest person I have ever known!!!